find dy/dx when
y= (x)^n + (n)^x + (x)^x + (n)^n
 

Grade:10

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
\\ \\\frac{d}{dx}\left(x^n+n^x+x^x+n^n\right)=nx^{n-1}+n^x\ln \left(n\right)+x^x\left(\ln \left(x\right)+1\right) \\
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty